# via Mixture-of-Curvature Experts

### v1

@@ 5.2 Ablating Distinct Curvature Learning with HELM-MiCE
[…]
HELM-MiCE 120M $\underline{19.3}$ $\mathbf{26.0}$ $\mathbf{27.4}$ $\mathbf{24.7}$ $\mathbf{23.5}$ $\mathbf{24.2}$
To assess the effectiveness of HELM-MiCE when each expert operates in a distinct curvature space, we additionally train a 120M-parameter HELM-MiCE model where the curvature of each expert is fixed to $-1.0$, which we denote as MiCE-Const. Consequently, MiCE-Const embeds the entire token sequence into a fixed space of curvature $-1.0$ similar to a dense model. MiCE-Const is trained with the same setup as the preceding models. We show the results in Table 2. HELM-MiCE outperforms the constant-curvature MiCE-Const in 4 out of the 5 benchmarks and achieves the higher overall accuracy, demonstrating the effectiveness of learning more expressive presentation by setting each expert to learn within a distinct manifold. Notably, MiCE-Const still outperforms the Euclidean DeepSeekV3 baseline on all 5 of the benchmarks, further demonstrating the effectiveness of hyperbolic LLMs over their Euclidean counterparts.
@@ 6 Conclusion
In this work, we introduce HELM, a family of fully hyperbolic large language models trained at hundred-million and billion-parameter scales. Operating entirely in hyperbolic space, HELM models are better aligned with the variable geometric structure of text and token distributions. We develop MiCE modules to construct HELM-MiCE variants, enabling fine-grained geometric learning and more expressive and geometrically flexible hidden representations. We further introduce HMLA mechanism to enable HELM models to be memory efficient and improve scalability. We also introduce the HoPE and $\text{{RMSNorm}}_{\mathcal{L}}$ modules, which are fundamental to building modern hyperbolic LLMs, and support them with extensive theoretical analysis and guarantees. Trained on 5B tokens, HELM models outperform their Euclidean counterparts across benchmarks in STEM reasoning, commonsense reasoning, and general knowledge. Nevertheless, the research presented has a few limitations. Due to computational constraints, our experiments only compare HELM to Euclidean LLMs trained on the same 5B tokens, which have less representational capacity when compared to the commercially available LLMs trained on much more extensive data [4, 1, 18, 42, 10]. Additionally, we chose the Wikipedia dataset for its widely accepted reliability. However, the trained models might be under-exposed to areas such as mathematical reasoning as a result. Future work could explore incorporating scaling laws [28, 25] for hyperbolic LLMs across larger compute and data frontiers to investigate their potential.
@@ Acknowledgments
This work was supported in part by the National Science Foundation (NSF) IIS Div Of Information & Intelligent Systems 2403317. We also gratefully acknowledge support in part from the Silicon Valley Community Foundation, an Amazon research award, the Yale AI Engineering Research Grant from Yale Office of the Provost, and an LEAP-U Sponsored Research from Samsung Research America. Additionally, this research has greatly benefited from the discussions and research talks held at the IMS-NTU Joint Workshop on Applied Geometry for Data Sciences. We also thank Ngoc Bui (Yale Univeristy) for useful feedback and discussion.
@@ A.1 Ollivier-Ricci Curvature
$\mu_{i}(j)$ =$=\begin{cases}\frac{1}{\operatorname{deg}(i)}&(i,j)\in E\\ 0&\text{otherwise.}\end{cases}$
@@
Here, $W^{\mathcal{G}}_{1}$ is the 1-Wasserstein distance between measures, $\mu_{i}$ and $\mu_{j}$. Intuitively, the curvature is defined as random walks between $i$ and $j$. If the random walks tend to stay at equal distances, $\kappa(i,j)=0$; if they diverge, $\kappa(i,j)<0$; and if they converge, $\kappa(i,j)>0$. Since curvature is a local property of a Riemannian surface, for graphs, we examine local neighborhoods with a step size of 1. We thus choose $\mu_{i}(j)=\frac{1}{\operatorname{deg}(j)}$ if $(i,j)\in E$, otherwise 00.
In our preliminary analysis of popular decoder-only LLMs (Figure 1), we draw $k$-nearest-neighbors graphs with the final-layer token embeddings for a collection of prompts from RedPajama [46] to ascertain their geometric structure. We observe a high variability of negative curvature values across tokens, hinting at their non-Euclidean nature. Robinson et al. [38] further allude to the non-Euclidean token subspace, necessitating language models that can accommodate this unique geometry.
@@ A.2 HoPE is a Lorentz Rotation
Let $\mathbf{X}$ be $T$ tokens with $\mathbf{x}_{i}\in\mathbb{L}^{K,d}$. Let $\mathbf{Q},\mathbf{K}$ be queries and keys as in Equation 3. Then $-d^{2}_{\mathcal{L}}\left(\mathrm{\mathrm{HoPE}\left(\mathbf{q}_{a}\right),\mathrm{HoPE}\left(\mathbf{k}_{b}\right))}\right)=g(\mathbf{x}_{a},\mathbf{x}_{b};a-b)$ for some function $g$.
@@
We will denote $\mathrm{HoPE}(\mathbf{q_{a}}),\mathrm{HoPE}(\mathbf{k_{b}})$ as $f_{q}(\mathbf{x}_{a}),f_{k}(\mathbf{x}_{b})$ where $f_{q},f_{k}$ denotes the function that projects the word embeddings to queries and keys, and then applying HoPE. In practice, the projection is done through a hyperbolic linear layer, which we take to be $\mathrm{HLT}$ from Section 2. It suffices to prove this proposition for the case of $d=2$, since HoPE does not affect the time-like dimension of the inputs and $\mathbf{R}_{i,\Theta}$ acts independently on each 2D block. First, note that we have
@@
[…]
@@
Now, given the fact that HoPE acts via Euclidean rotation on the time-like dimension of any vector, we have $\varphi_{q}(\mathbf{x}_{a},a)=\varphi_{q}(\mathbf{x}_{a},a^{\prime}),\varphi_{k}(\mathbf{x}_{b},b)=\varphi_{k}(\mathbf{x}_{b},b^{\prime})$ for any $a^{\prime},b^{\prime}$. In particular, when $a^{\prime}=b^{\prime}=0$, HoPE acts via identity since all rotation angles become 00. Hence, we have
@@
Furthermore, we have $\varphi_{g^{\prime}}=\|(\mathbf{q}_{a})_{s}\|\|(\mathbf{k}_{b})_{s}\|=\|\mathbf{W^{Q}}\mathbf{x}_{a}\|\|\mathbf{W^{K}}\mathbf{x}_{b}\|$ by plugging back into Equation 9, which is a function of just the word embeddings. Next, for the angular component, note that given the definition of HoPE, the rotation on any 2D block of the space-like dimension of the input at position $p$ is simply a scaling of a fixed rotation angle by $p$. Letting this fixed angle be $\sigma$, the rotation is precisely $p\sigma$. Therefore, we have
@@
where $\theta_{q},\theta_{k}$ denote the angular components of $\mathbf{q}_{a},\mathbf{k}_{b}$. Next, given Equation 9, we have $\vartheta_{g^{\prime}}=(a-b)(\sigma)+(\theta_{q}-\theta_{k})$. Note that we have $e^{i\theta_{q}}=\frac{\mathbf{W^{Q}}\mathbf{x}_{a}}{\|\mathbf{W^{Q}}\mathbf{x}_{a}\|}$ and $e^{i\theta_{k}}=\frac{\mathbf{W^{K}}\mathbf{x}_{b}}{\|\mathbf{W^{K}}\mathbf{x}_{b}\|}$. Consequently, $\vartheta_{g^{\prime}}$ is a function of the word embeddings and the relative position $a-b$. All in all, $-d^{2}_{\mathcal{L}}(f_{q}(\mathbf{x}_{a}),f_{k}(\mathbf{x}_{b}))$ can be expressed with a function $g(\mathbf{x}_{a},\mathbf{x}_{b};a-b)$ as desired. ∎
@@
Let $\mathbf{Q},\mathbf{K}$ be as defined in Equation 3, then the negative square Lorentz distance $-d_{\mathcal{L}}\left(\mathrm{HoPE}\left(\mathbf{q}_{a}\right),\mathrm{HoPE}(\mathbf{k}_{b})\right)$ can be upper bounded by $f(\mathbf{q}_{a},\mathbf{k}_{b})g(a-b)<0$, where $f$ has no dependencies on position, and $g$ depends entirely on relative position and scales inversely w.r.t. $a-b$.
@@
$-d_{\mathcal{L}}^{2}\left(\mathrm{HoPE}\left(\mathbf{q}\right),\mathrm{HoPE}(\mathbf{k})\right)=-\frac{2}{K}-2(q_{t}k_{t})+2\mathbf{q}_{s}^{\top}\mathbf{k}_{s}.$
Next, for simplicity, we denote $\mathbf{q}_{s},\mathbf{k}_{s}$ as $\mathbf{a},\mathbf{b}$ respectively. We group together entries of the queries and keys, where $\mathbf{a}_{[2k:2k+1]},\mathbf{b}_{[2k:2k+1]}$ as the $2k$-th and the $(2k+1)$-th entries of $\mathbf{a},\mathbf{b}$ respectively. With this in mind, note that since we take query and key projects to be with $\mathrm{HLT}$ as given in Section 2, we obtain $\mathbf{a}=\mathbf{W^{Q}}\mathbf{x}_{a}$ and $\mathbf{b}=\mathbf{W^{K}}\mathbf{x}_{b}$. To that end, we can assert that
@@
Now, note that the term $\mathbf{A}_{k+1}-\mathbf{A}_{k}$ has no dependency on position. The sum $\sum_{k=0}^{n/2}|\mathbf{E}_{k}|$ scales inversely with $b-a$, as shown by Su et al. [40]. To that end, we have
@@
Let $\mathbf{Q},\mathbf{K}$ be as defined in Equation 3, then HoPE can be maximal at an arbitrary distance, i.e., for any relative distance $r\in\mathbb{Z}$, there exists a key $\mathbf{k}_{j}$ such that the softmax value is maximum at distance $r$.
@@
We first restate Lemma A.1. from Barbero et al. [2]. The remainder of our proof follows a similar layout to that of Proposition 3.1. in Barbero et al. [2].
Lemma A.1. (Barbero et al. [2])    Consider $g\in\mathbb{Q}$ with $g\neq 0$ and $n\in\mathbb{Z}$. Then, $ng\equiv 0~{}(\operatorname{mod}~{}2\pi)$ only when $n=0$. In particular, this also holds if $g$ is algebraic.
Consider a distance $r$, a non-trivial query $\mathbf{Q}_{p}=\psi\in\mathbb{L}^{K,n}$, as well as a key $\mathbf{K}=\operatorname{HoPE}(\mathbf{Q}_{p})$. We can represent $\psi$ as a combination of a time and space dimension on the Lorentzian manifold, $[\psi_{t},\psi_{s}]\in\mathbb{L}^{K,n}$. Using our definition of HoPE in Section 4.1, we have
@@
Using Lemma A.1. from Barbero et al. [2], we observe the maximum can be achieved when $j-i=-r$ for $j-i\leq 0$ since we are using causal masking, $j\leq i$. This ensures $\cos{(j-i+r)\theta_{l}}=\cos{(0)}=1$, concluding the proof. ∎
@@
The proof follows a similar layout as that of Proposition 5.3. from Barbero et al. [2]. We start with the diagonal case. Suppose $\mathbf{Q}_{i}=\mathbf{K}_{j}=\psi\in\mathbb{L}^{K,d}$, for non-trivial $\psi=[\psi_{t},\psi_{s}]$. We assume embedding dimension $d=2$, i.e., only a single rotation block $\mathbf{R}_{i,\theta}$ acts on the embeddings.
@@
Using Lemma A.1. from [2], when $j=i$, we have $(j-i)\theta\equiv 0~{}(\text{mod}~{}2\pi)$. Next, let us define $\mathbf{a}_{i,i}$ as $-[C-2\|\psi_{s}\|^{2}]$. This means $\cos{((j-i)\theta)}=\cos{(0)}=1$, and $\mathbf{a}_{i,j}<\mathbf{a}_{i,i}$. This gives us the following self-attention score:
@@
for all $k\neq i$, $\cos{((k-i)\theta)}<1$. This means $02\|\psi_{s}\|^{2}(\cos{((k-i)\theta)}-1)<$ 0. To that end, we obtain
@@
[…]
@@
where $\mathbf{v}$ denotes the inputs to the activation $\sigma$. These gradients are invariant to the scaling of Euclidean inputs $x$ and weights $\mathbf{W}$, trivially for $\frac{\partial L}{\partial\mathbf{b}}$, and due to the linearity established in Equation (14) for $\frac{\partial L}{\partial\mathbf{g}}$. Computing $\frac{\partial L}{\partial\mathbf{W}}$ is more involved due to the quadratic computation in $\operatorname{RMS}$, but also provides invariance to input scaling as shown by Zhang and Sennrich [51].

### v2

@@ 5.2 Ablating Distinct Curvature Learning with HELM-MiCE
MiCE-Const 120M $\mathbf{20.0}$ $\underline{25.6}$ $\underline{27.0}$ $23.5$ $\underline{22.3}$ $\underline{23.7}$
HELM-MiCE 120M $\underline{19.7\pm 0.3}$ $\mathbf{25.9}\pm 0.2$} $\mathbf{27.7}\pm 0.4$ $\mathbf{24.4}\pm 0.2$ $\mathbf{23.2}\pm 0.5$ $\mathbf{24.1}\pm 0.1$
To assess the effectiveness of each expert operating in a distinct curvature space, we train a 120M-parameter HELM-MiCE model where the curvature of each expert is fixed to 1.0-1.0, which we denote as MiCE-Const. Consequently, MiCE-Const embeds the entire token sequence into a fixed space of curvature 1.0-1.0 similar to a dense model. MiCE-Const is trained with the same setup as the preceding models. We show the results in Table˜2. HELM-MiCE outperforms the constant-curvature MiCE-Const in 4 out of the 5 benchmarks and achieves the higher overall accuracy, demonstrating the effectiveness of learning more expressive presentation by setting each expert to learn within a distinct manifold. Notably, MiCE-Const still outperforms the Euclidean DeepSeekV3 baseline on all 5 of the benchmarks, further demonstrating the effectiveness of hyperbolic LLMs over their Euclidean counterparts. We don’t provide statistics for ablation baselines due to the need for excessive compute. Nevertheless, the results remain statistically significant as demonstrated in Table˜2.
5.3 Qualitative Studies on Semantic Hierarchy Modeling
In this section, we qualitatively access the ability of HELM to model semantic hierarchy against its Euclidean counterparts. Our investigation of final-layer embedding distributions have found that HELM learns representations where more generic words tend to cluster in areas of smaller norm and more specific words tend to have larger norms. In Table˜3, we provide case studies for HELM-MiCE 1B and DeepseekV3 1B, where we show embedding norm in the final layers for words of varying levels of specificity (top table) and for a sample question taken from the MMLU benchmark. For HELM-MiCE, more generic words (e.g., subject) are clustered closer to the origin than more specific words (e.g., biology), which has a smaller norm than even more specific words (e.g., photosynthesis). However, this does not necessarily hold for the DeepseekV3 1B model, demonstrating how HELM-MiCE better handles semantic hierarchies. We will provide low-dimensional visualization of these embeddings in our revision. The hierarchical organization of the space enables the HELM models to sometimes better navigate the embedding space and obtain better performance.
@@ 6 Conclusion
In this work, we introduce HELM, a family of fully hyperbolic large language models trained at hundred-million and billion-parameter scales. Operating entirely in hyperbolic space, HELM models are better aligned with the variable geometric structure of text and token distributions. We develop MiCE modules to construct HELM-MiCE variants, enabling fine-grained geometric learning and more expressive and geometrically flexible hidden representations. We further introduce HMLA mechanism to enable HELM models to be memory efficient and improve scalability. We also introduce the HoPE and $\text{{RMSNorm}}_{\mathcal{L}}$ modules, which are fundamental to building modern hyperbolic LLMs, and support them with extensive theoretical analysis and guarantees. Trained on 5B tokens, HELM models outperform their Euclidean counterparts across benchmarks in STEM reasoning, commonsense reasoning, and general knowledge. Nevertheless, the research presented has a few limitations. Due to computational constraints, our experiments only compare HELM to Euclidean LLMs trained on the same 5B tokens, which have less representational capacity when compared to the commercially available LLMs trained on much more extensive data [gpt-neox-20b, achiam2023gpt, dubey2024llama3, gemma2024, deepseekai2024deepseekv3technicalreport]. Additionally, we chose the Wikipedia dataset for its widely accepted reliability. However, the trained models might be under-exposed to areas such as mathematical reasoning as a result. Future work could explore incorporating scaling laws [kaplan2020scalinglawsneurallanguage, hoffmann2022trainingcomputeoptimallargelanguage] for hyperbolic LLMs across larger compute and data frontiers to investigate their potential.
Table 3: Case study investigate of embedding norm in the final layer of HELM (1B) and DeepseekV3 (1B). Top: embedding norm of words of varying levels of specificity; Bottom: embedding norm of a question taken from the MMLU benchmark. For HELM-MiCE, more generic words are clustered closer to the origin than more specific words, which has a smaller norm than even more specific words. However, this does not necessarily hold for the DeepseekV3 1B model.
Words HELM-MiCE DeepseekV3
Average Norm Range Average Norm Range
to, in, have, that, and, is, for 35.930 35.890~35.951 33.725 33.660~33.800
study, research, subject, papers, category 36.080 36.030~36.033 33.735 33.668~33.776
biology, physics, chemistry, mathematics, computer science 36.155 36.033~36.270 33.720 33.658~33.776
algebra, geometry, photosynthesis, cellular respiration, genetics 36.288 36.133~36.484 33.741 33.622~33.826
HELM-MiCE DeepseekV3
Words Norm Range Words Norm Range
A, How, does, if, there, have, is, any, with, of 36.031~36.396 is, a, connecting, graph, there, edges, complete, have, of 33.668~33.768
discrete, vertices, edges, connecting, pair, graph, complete, many, 10 36.506~36.717 discrete, 10, how, if, pair, does, with, A, vertices, any 33.772~33.908
@@ Acknowledgments
This work was supported in part by the National Science Foundation (NSF) IIS Div Of Information & Intelligent Systems 2403317 and Army Research Office contract W911NF-23-1-0088. We also acknowledge support in part from the Silicon Valley Community Foundation, an Amazon research award, the Yale AI Engineering Research Grant from Yale Office of the Provost, and an LEAP-U Sponsored Research from Samsung Research America. Moreover, this research has greatly benefited from the discussions and research talks held at the IMS-NTU Joint Workshop on Applied Geometry for Data Sciences. We also thank Ngoc Bui (Yale Univeristy) for useful feedback and discussion.
@@ A.1 Ollivier-Ricci Curvature
$\mu_{i}(j)$ $=\begin{cases}\frac{1}{\operatorname{deg}(i)}&(i,j)\in E\\ 0&\text{otherwise.}\end{cases}$
@@
Here, $W^{\mathcal{G}}_{1}$ is the 1-Wasserstein distance between measures, $\mu_{i}$ and $\mu_{j}$. Intuitively, the curvature is defined as random walks between $i$ and $j$. If the random walks tend to stay at equal distances, $\kappa(i,j)=0$; if they diverge, $\kappa(i,j)<0$; and if they converge, $\kappa(i,j)>0$. Since curvature is a local property of a Riemannian surface, for graphs, we examine local neighborhoods with a step size of 1. We thus choose $\mu_{i}(j)=\frac{1}{\operatorname{deg}(j)}$ if $(i,j)\in E$, otherwise 0.
In our preliminary analysis of popular decoder-only LLMs (Figure 1), we draw $k$-nearest-neighbors graphs with the final-layer token embeddings for a collection of prompts from RedPajama [weber2024redpajamaopendatasettraining] to ascertain their geometric structure. We observe a high variability of negative curvature values across tokens, hinting at their non-Euclidean nature. robinson2024structuretokenspacelarge further allude to the non-Euclidean token subspace, necessitating language models that can accommodate this unique geometry.
@@ A.2 HoPE is a Lorentz Rotation
Let $\mathbf{X}$ be $T$ tokens with $\mathbf{x}_{i}\in\mathbb{L}^{K,d}$. Let $\mathbf{Q},\mathbf{K}$ be queries and keys as in Equation˜3. Then $d^{2}_{\mathcal{L}}\left(\mathrm{\mathrm{HoPE}\left(\mathbf{q}_{a}\right),\mathrm{HoPE}\left(\mathbf{k}_{b}\right))}\right)=g(\mathbf{x}_{a},\mathbf{x}_{b};a-b)$ for some function $g$.
@@
We will denote $\mathrm{HoPE}(\mathbf{q_{a}}),\mathrm{HoPE}(\mathbf{k_{b}})$ as $f_{q}(\mathbf{x}_{a}),f_{k}(\mathbf{x}_{b})$ where $f_{q},f_{k}$ denotes the function that projects the word embeddings to queries and keys, and then applying HoPE. In practice, the projection is done through a hyperbolic linear layer, which we take to be $\mathrm{HLT}$ from Section˜2. It suffices to prove this proposition for the case of $d=2$, since HoPE does not affect the time-like dimension of the inputs and $\mathbf{R}_{i,\Theta}$ acts independently on each 2D block. First, note that we have
@@
where $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ denotes Lorentzian inner product and $\langle\cdot,\cdot\rangle$ denotes the regular Euclidean inner product. Since HoPE is a Lorentz rotation, the term $2\left(f_{q}(\mathbf{x}_{a})_{t}f_{k}(\mathbf{x}_{b})_{t}\right)$ is simply $2((\mathbf{q}_{a})_{t},(\mathbf{k}_{b})_{t})=2\left(\sqrt{\mathbf{W^{Q}}\mathbf{x}_{a}-1/K}\sqrt{\mathbf{W^{K}}\mathbf{x}_{b}-1/K}\right)$, so we focus on the inner product $\langle f_{q}(\mathbf{x}_{a})_{s},f_{k}(\mathbf{x}_{b})_{s}\rangle$. Then, by assuming that $d=2$, we have $(f_{q}(\mathbf{x}_{a}))_{s},(f_{k}(\mathbf{x}_{b}))_{s}\in\mathbb{R}^{2}$; hence, we can parametrize these vectors by their radial and angular components. For simplicity, denote $(f_{q}(\mathbf{x}_{a}))_{s},(f_{k}(\mathbf{x}_{b}))_{s}$ as $\mathbf{a},\mathbf{b}$ respectively. Then, write $\langle\mathbf{a},\mathbf{b}\rangle$ as a function $g^{\prime}$. Afterwards, we parametrize the vectors as
@@
Now, given the fact that HoPE acts via Euclidean rotation on the time-like dimension of any vector, we have $\varphi_{q}(\mathbf{x}_{a},a)=\varphi_{q}(\mathbf{x}_{a},a^{\prime}),\varphi_{k}(\mathbf{x}_{b},b)=\varphi_{k}(\mathbf{x}_{b},b^{\prime})$ for any $a^{\prime},b^{\prime}$. In particular, when $a^{\prime}=b^{\prime}=0$, HoPE acts via identity since all rotation angles become 0. Hence, we have
@@
Furthermore, we have $\varphi_{g^{\prime}}=\|(\mathbf{q}_{a})_{s}\|\|(\mathbf{k}_{b})_{s}\|=\|\mathbf{W^{Q}}\mathbf{x}_{a}\|\|\mathbf{W^{K}}\mathbf{x}_{b}\|$ by plugging back into Equation˜9, which is a function of just the word embeddings. Next, for the angular component, note that given the definition of HoPE, the rotation on any 2D block of the space-like dimension of the input at position $p$ is simply a scaling of a fixed rotation angle by $p$. Letting this fixed angle be $\sigma$, the rotation is precisely $p\sigma$. Therefore, we have
@@
where $\theta_{q},\theta_{k}$ denote the angular components of $\mathbf{q}_{a},\mathbf{k}_{b}$. Next, given Equation˜9, we have $\vartheta_{g^{\prime}}=(a-b)(\sigma)+(\theta_{q}-\theta_{k})$. Note that we have $e^{i\theta_{q}}=\frac{\mathbf{W^{Q}}\mathbf{x}_{a}}{\|\mathbf{W^{Q}}\mathbf{x}_{a}\|}$ and $e^{i\theta_{k}}=\frac{\mathbf{W^{K}}\mathbf{x}_{b}}{\|\mathbf{W^{K}}\mathbf{x}_{b}\|}$. Consequently, $\vartheta_{g^{\prime}}$ is a function of the word embeddings and the relative position aba-b. All in all, $d^{2}_{\mathcal{L}}(f_{q}(\mathbf{x}_{a}),f_{k}(\mathbf{x}_{b}))$ can be expressed with a function $g(\mathbf{x}_{a},\mathbf{x}_{b};a-b)$ as desired. ∎
@@
Let $\mathbf{Q},\mathbf{K}$ be as defined in Equation˜3, then the negative square Lorentz distance $d_{\mathcal{L}}\left(\mathrm{HoPE}\left(\mathbf{q}_{a}\right),\mathrm{HoPE}(\mathbf{k}_{b})\right)$ can be upper bounded by $f(\mathbf{q}_{a},\mathbf{k}_{b})g(a-b)<0$, where $f$ has no dependencies on position, and $g$ depends entirely on relative position and scales inversely w.r.t. aba-b.
@@
$d_{\mathcal{L}}^{2}\left(\mathrm{HoPE}\left(\mathbf{q}\right),\mathrm{HoPE}(\mathbf{k})\right)=-\frac{2}{K}-2(q_{t}k_{t})+2\mathbf{q}_{s}^{\top}\mathbf{k}_{s}.$
Next, for simplicity, we denote $\mathbf{q}_{s},\mathbf{k}_{s}$ as $\mathbf{a},\mathbf{b}$ respectively. We group together entries of the queries and keys, where $\mathbf{a}_{[2k:2k+1]},\mathbf{b}_{[2k:2k+1]}$ as the $2k$-th and the $(2k+1)$-th entries of $\mathbf{a},\mathbf{b}$ respectively. With this in mind, note that since we take query and key projects to be with $\mathrm{HLT}$ as given in Section˜2, we obtain $\mathbf{a}=\mathbf{W^{Q}}\mathbf{x}_{a}$ and $\mathbf{b}=\mathbf{W^{K}}\mathbf{x}_{b}$. To that end, we can assert that
@@
Now, note that the term $\mathbf{A}_{k+1}-\mathbf{A}_{k}$ has no dependency on position. The sum $\sum_{k=0}^{n/2}|\mathbf{E}_{k}|$ scales inversely with bab-a, as shown by su2021roformer. To that end, we have
@@
Let $\mathbf{Q},\mathbf{K}$ be as defined in Equation˜3, then HoPE can be maximal at an arbitrary distance, i.e., for any relative distance $r\in\mathbb{Z}$, there exists a key $\mathbf{k}_{j}$ such that the softmax value is maximum at distance $r$.
@@
We first restate Lemma A.1. from barbero2025rope. The remainder of our proof follows a similar layout to that of Proposition 3.1. in barbero2025rope.
Lemma A.1. (barbero2025rope)    Consider $g\in\mathbb{Q}$ with $g\neq 0$ and $n\in\mathbb{Z}$. Then, $ng\equiv 0~(\operatorname{mod}~2\pi)$ only when $n=0$. In particular, this also holds if $g$ is algebraic.
Consider a distance $r$, a non-trivial query $\mathbf{Q}_{p}=\psi\in\mathbb{L}^{K,n}$, as well as a key $\mathbf{K}=\operatorname{HoPE}(\mathbf{Q}_{p})$. We can represent $\psi$ as a combination of a time and space dimension on the Lorentzian manifold, $\psi_{t},\psi_{s}]\in\mathbb{L}^{K,n}$. Using our definition of HoPE in Section 4.1, we have
@@
Using Lemma A.1. from barbero2025rope, we observe the maximum can be achieved when ji=rj-i=-r for $j-i\leq 0$ since we are using causal masking, $j\leq i$. This ensures $\cos{(j-i+r)\theta_{l}}=\cos{(0)}=1$, concluding the proof. ∎
@@
The proof follows a similar layout as that of Proposition 5.3. from barbero2025rope. We start with the diagonal case. Suppose $\mathbf{Q}_{i}=\mathbf{K}_{j}=\psi\in\mathbb{L}^{K,d}$, for non-trivial $\psi=[\psi_{t},\psi_{s}]$. We assume embedding dimension $d=2$, i.e., only a single rotation block $\mathbf{R}_{i,\theta}$ acts on the embeddings.
@@
Using Lemma A.1. from [barbero2025rope], when $j=i$, we have $(j-i)\theta\equiv 0~(\text{mod}~2\pi)$. Next, let us define $\mathbf{a}_{i,i}$ as $C-2\|\psi_{s}\|^{2}]$. This means $\cos{((j-i)\theta)}=\cos{(0)}=1$, and $\mathbf{a}_{i,j}<\mathbf{a}_{i,i}$. This gives us the following self-attention score:
@@
for all $k\neq i$, $\cos{((k-i)\theta)}<1$. This means $2\|\psi_{s}\|^{2}(\cos{((k-i)\theta)}-1)<0$. To that end, we obtain
@@
$=\sigma\Bigg(\frac{\cancel{\delta}\mathbf{W}^{\top}\mathbf{x}}{\cancel{\delta}\mathrm{RMS}(\mathbf{W}^{\top}\mathbf{x})}\odot\mathbf{g}+\mathbf{b}\Bigg)=\mathbf{y}$ (14)
@@
where $\mathbf{v}$ denotes the inputs to the activation $\sigma$. These gradients are invariant to the scaling of Euclidean inputs $x$ and weights $\mathbf{W}$, trivially for $\frac{\partial L}{\partial\mathbf{b}}$, and due to the linearity established in Equation (14) for $\frac{\partial L}{\partial\mathbf{g}}$. Computing $\frac{\partial L}{\partial\mathbf{W}}$ is more involved due to the quadratic computation in $\operatorname{RMS}$, but also provides invariance to input scaling as shown by zhang2019rootmeansquarelayer.